Can someone please help me on this three questions
Pre-Lab Questions Ansuer the follow ingquestams before comm๕ to class 1. The following data were obtained in an experiment to determine the enthalpy of fusion of water. Calculate the enthalpy of fusion of ice (in J/g) from the data. Asume the specific heat of liquid water is 4.1841g。 Mass of water in calorimeter = 108.31 g Initial temperature of water 25.1C Mass of ice at 0°C #5.35g Final temperature of the water with melted ice = 19.6 If a solution is 0.84% oleic acid and 107 drops of the solution have a total volume of 13 mL, calculate the volume of pure oleic acid in one drop of solution. 2. 3. Sketch a body-centered cubic unit cell and a face-centered cubic unit cell. CH
Answer #1

Solution :-

Q1) Amount of heat loss by the water will be absorbed by the ice which causes the melting of ice and then melted ice water raises in the temperature from 0 degree to 19.6 degree C.

So we can write it as

-q = +q

-m*c*delta T = (m ice *delta H fus ice)+(m*c*delta T)

-108.31 g* 4.184 J/gC*(19.6C -25.1C) = (5.35g* delta H fus ice)+(5.35g*4.184J/gC*(19.6 C – 0.0C))

2492.43 J = (5.35g* delta H fus ice) + 438.7 J

2492.43 J – 438.7 J = (5.35g* delta H fus ice)

2053.3 J = (5.35g* delta H fus ice)

2053.3 J / 5.35 g = delta H fus

384 J/g = delta H fus

Therefore the enthalpy of fusion of ice is 384 J/g